OpenStudy (a1234):

Is it possible to simplify a radical expression like this? 60(√3/36) Dividing 36 and 60 by 2 to simplify: 5(√5/3) Is this fine?

OpenStudy (whpalmer4):

Yes, that is.

OpenStudy (a1234):

So, it is fine to simplify that way?

OpenStudy (whpalmer4):

Well, let me qualify that: if you do it right, it is :-) \[60(\frac{\sqrt{3}}{36}) = (\frac{60}{36})\sqrt{3} = (\cancel{\frac{12}{12}}*\frac{5}{3})\sqrt{3} = 5(\frac{\sqrt{3}}{3}) \] I assume you just made a typo...

OpenStudy (a1234):

That's what I mean :). Thanks!